### v1

# SonicMotion: Dynamic Spatial Audio Soundscapes with Latent Diffusion Models

Abstract
Spatial audio is an integral part of immersive entertainment, such as VR/AR, and has seen increasing popularity in cinema and music as well. The most common format of spatial audio is described as first-order Ambisonics (FOA). We seek to extend recent advancements in FOA generative AI models to enable the generation of 3D scenes with dynamic sound sources. Our proposed end-to-end model, SonicMotion, comes in two variations which vary in their user input and level of precision in sound source localization. In addition to our model, we also present a new dataset of simulated spatial audio-caption pairs. Evaluation of our models demonstrate that they are capable of matching the semantic alignment and audio quality of state of the art models while capturing the desired spatial attributes.
Index Terms:
Spatial audio generation, Ambisonic audio generation, Latent diffusionI Introduction
The increasing use of spatial audio formats for immersive entertainment has created an opportunity for generative machine learning to simplify the production of first-order Ambisonics (FOA) soundscapes. The 4-channel format of FOA audio provides precise control over the direction of arrival of sound sources in the X, Y, and Z directions. It is commonly implemented in applications including VR/AR and music, where the directionality of a sound source has a significant impact on immersion. Creating spatial audio for these purposes can be time consuming, as it requires manually panning sources to the desired location using specialized software. An opportunity to ease this problem through natural language descriptions and audio generation models presents itself. In recent years, generative audio models have seen significant advances in speech, sound clip, and music generation. However, most of these models are limited to generating mono or stereo audio and do not consider spatial locations of audio sources or environmental conditions [1, 2, 3].
Several models have since been proposed to enable generation of audio with the ability to control spatial locations and environmental conditions. In [4], the authors used the diffusion transformer model from [3] to develop SpatialSonic—a stereo audio generation model with spatial location awareness. The model is conditioned using a spatial text prompt, as well as several integer parameters which represent the location of the sound source. The user can therefore control the initial position, final position, and speed of the sound source. However, as the model only generates stereo audio, it is limited to spatial cues such as “left,” “front,” and “right” and cannot capture cues such as “behind,” or “above.”
In [5], the authors propose Diff-SAGe, a flow-based diffusion transformer, to enable the generation of FOA audio. The model is conditioned using a unified representation of a class embedding, an azimuthal angle embedding, and an elevation angle embedding. While this model is capable of generating FOA audio with precise azimuth and elevation control, the position of the source must remain static. Another model which generates FOA audio, ImmerseDiffusion [6], utilizes the same diffusion transformer as [3] with some adaptations for FOA audio. Additionally, the model is conditioned solely on a text prompt using the ELSA [7] text encoder, which enables spatially aware text and audio embedding. This model is successful in generating FOA audio with spatial location guidance via natural language, however it does not support moving sound sources.
In this paper we propose SonicMotion, a new generative spatial audio model which generates FOA audio with support for precise control over moving sound sources. Our model comes in two variations. The first variation is conditioned using text, while the second variation is conditioned using text and spatial parameters for a greater level of precision. In addition to our model, we also present a new dataset of spatial audio and spatial caption pairs. Our dataset includes samples with moving and static sources, along with recorded parameters such as directional angles and movement times. We evaluate our models using established metrics for fair comparison against state of the art models. Additional metrics are included to evaluate the spatial accuracy of our models.
II Methodology
The main components of our work include the dataset and the model, SonicMotion. The model includes three sub-components which are the autoencoder pretransform, conditioning, and the diffusion transformer. We additionally describe our training setup and the metrics used to evaluate both the autoencoder and overall end-to-end model.
II-A FOA Format
To give insight into the relevant audio format, we define Ambisonics as a 4-channel surround sound format which encodes sound direction to deliver a 3D audio experience. First-order Ambisonics (FOA) is the most commonly used format of Ambisonics due to its simplicity while maintaining sufficient spatial resolution. FOA represents spatial audio using four channels: W, X, Y, and Z. The W channel is omnidirectional and the rest of the channels represent the directional components of their respective cardinal planes. In equation set (1), represents the sound pressure i.e. a mono audio signal. and represent the azimuthal angle and angle of elevation respectively.
II-B Dataset
To our knowledge there are no publicly available datasets consisting of spatial audio-caption pairs which include samples with moving sources. To overcome this lack of data, we introduce a dataset which involves spatially augmenting three existing audio-caption pair datasets. We select Clotho [8], AudioCaps [9], and FreeSound [10] as original datasets to be spatially augmented. First, we resample the audio to 16khz, remove leading and trailing silences, and either truncate or loop the audio so that it is 10 seconds long. For each original audio sample, we generate one static and one dynamic spatial audio sample. To convert the mono audio signal to spatial audio we first randomly generate the initial azimuth and elevation angles to determine the direction. For spatial samples with a static source we can use these initial azimuth and elevation angles to directly encode the mono signal into FOA using equation set (1). For spatial samples which include a moving source we then randomly determine whether the azimuth, elevation, or both change over time. In any case, we randomly determine the final position of the angle such that it differs by at least for azimuth and for elevation. The speed of the moving source is randomly determined as “fast,” “moderate,” or “slow.” Movement start and end times are randomly determined based on the speed category. Finally, we linearly interpolate between initial and final angles in the movement window and encode into FOA using (1). When interpolating between initial and azimuth angles, we randomly determine if the movement is clockwise or counter-clockwise. For each sample we record the initial and final azimuth and elevation angles, a clockwise flag, speed, and movement start and end times.
Our dataset additionally includes a new set of captions which capture relevant details about the direction and movement of the sound source. To achieve this, we leverage Google’s Gemma-3 12B [11] to automatically rephrase the original caption when provided with a list of relevant spatial parameters. Spatial parameters and their mappings to natural language are detailed in Table I. The final dataset consists of a training split with 1,018,430 samples, a validation split with 3,078 samples and a test split with 4,040 samples.
| Feature | Range | Bins | Map |
| Azimuth | front front-left left back-left back back back-right right front-right | ||
| Elevation | down up | ||
| Speed | (1.0s, 10.0s) | (1.0s, 3.0s) (3.5s, 6.5s) (7.0s, 10.0s) | fast moderate slow |
II-C Spatial Autoencoder
We use an autoencoder as a pretransform for compressing audio training data into latent space and decoding generated latents back into the original representation space. Autoencoders are frequently used to fulfill this need as their encoder and decoder can be decoupled to support both training and inference. We closely follow [6] and use the autoencoder provided by Descript Audio Codec (DAC) [12] with some adaptations for FOA format. DAC is structured as a 1-dimensional U-Net autoencoder with a series of convolutions that repeatedly down-sample the input sequence during encoding and up-sample the encoded latents during decoding. To adapt DAC to our use case, we first replace the Residual Vector Quantization bottleneck with a continuous Variational Autoencoder bottleneck to avoid loss of information due to quantization. We remove the final tanh() activation from the decoder to prevent harmonic distortion [3]. Next, we define a new loss function which better represents the FOA format. We use the loss function in equation (2), which sums the Multi-Resolution Short Time Fourier Transform (MRSTFT) loss of each channel, the KL divergence loss of the bottleneck, as well as adversarial and feature matching losses applied to the discriminator. The parameter represents a weight which is perceptually updated during training, while the parameter represents a weight which remains constant. The autoencoder compresses a 4-channel FOA signal into 64 channels with a down-sampling ratio of 2048 and an overall compression factor of 128.
II-D Conditioning
We propose two models which differ only in the extent of their conditioning. The descriptive model is conditioned on a text prompt embedding while the parametric model is conditioned on the same text embedding as well as a positional embedding which represents the azimuth and elevation of the sound at any time. We encode the text prompt using the pre-trained T5 encoder [13]. As shown in an ablation study in [4], the spatial retrieval performance of the T5 encoder exceeds that of the CLAP [14] text encoder. To enhance the precision in the parametric model, we introduce an additional conditioner based on a state matrix that captures the position of the sound source with respect to time. We modify the position encoder from [4] which provides guidance for azimuth based on the initial position, final position, and speed of the sound source. Modifications are made to support the FOA domain. Namely, we expand the range of azimuth to span a full 360∘ and account for both clockwise and counter-clockwise movement. We also accept additional parameters to encode the initial and final elevation. To build the state matrix, we first use (3) to describe the position of the sound, , given a current time, , and duration, . We build a zeros matrix of shape where is the number of bins representing unique positions and assign a value of 1 to represent the current position of the sound according to (4). The guidance matrix is generated for both azimuth and elevation and these two matrices are concatenated. The resulting matrix is fused with the embedding of the text prompt using a cross-attention module with 4 layers and 4 heads. Our model also contains two additional cross-attention conditioners to include the temporal conditions, specifically start time and total time.
II-E Diffusion Transformer
We use the continuous diffusion transformer from [3]. The model includes blocks with self-attention, cross-attention and multi-layer perceptron gates with layer normalization and skip connections. The conditioning token dimension is set to 768 and operates with a projection layer to fit all of the conditioners. We train with the objective of minimizing the Mean Squared Error between the ground truth latent and the predicted latent at the given timestep, for the given conditions.
II-F Training
The autoencoder is trained using randomly cropped 2 second segments of audio from the training split of the spatial dataset on a cluster of 4xH100 GPUs. We use a batch size of 128 and train for 400K steps. The autoencoder generator uses an initial learning rate of , while the discriminator uses an initial learning rate of . Both the generator and discriminator use the AdamW optimizer with momentum values of 0.8 and 0.99 and a weight decay factor of .
The diffusion model is similarly trained on a cluster of 4xH100 GPUs. We use the full 10 seconds of audio from each sample in the training set. During training, we freeze the weights of the autoencoder and pre-trained T5 text encoder. We note that the weights of the cross attention mechanism in the position encoder remain unfrozen and are learned during training of the diffusion model. The model is trained with a batch size of 1024 for 100K steps. We use the AdamW optimizer with a learning rate of , momentum values of 0.9 and 0.99 and a weight decay factor of . The cross attention mechanism of the position encoder uses an increased learning rate of We additionally employ an inverse learning rate scheduler with an inverse gamma of , power of 0.5 and warm up of 0.99. The descriptive model is trained using only spatial captions while the parametric model is trained for 50k steps using the original non-spatial captions and 50k steps using the spatial captions.
| Model | CLAP | FD | FAD | KL | L1() | L1() | |
| AudioLDM 2 | 0.22 | 0.55 | 1.78 | 0.27 | - | - | - |
| Stable Audio Open | 0.19 | 1.18 | 7.90 | 0.59 | - | - | - |
| SonicMotion (Descriptive) | 0.23 | 0.52 | 2.97 | 0.26 | 21.20∘ | 16.04∘ | 29.22∘ |
| SonicMotion (Parametric) | 0.23 | 0.57 | 3.00 | 0.28 | 13.17∘ | 4.01∘ | 14.32∘ |
II-G Evaluation
We evaluate the reconstruction quality of the autoencoder on the STFT and Mel Spectrogram (MEL) distance between the original audio and its reconstructed counterpart. These evaluation metrics are supported by the AuraLoss [15] library and are implemented using the default settings. To evaluate the spatial accuracy of reconstructed audio we derive the intensity vectors for each cardinal direction by multiplying the omnidirectional W channel with the X, Y, or Z channel as shown in (5). We then calculate the vectors representing azimuth () and elevation () at each sample in the audio using (6). The error is computed as the L1 norm of the circular difference in azimuth and the linear difference in elevation between the original and reconstructed audio [6].
We additionally report the error in direction of arrival (DoA) between original and reconstructed audio. This difference in spatial angle is calculated with (7) and (8) below:
The diffusion model is evaluated using CLAP score, Fréchet Distance (FD), Fréchet Audio Distance (FAD), and Kullback-Leibler Divergence (KL). To compute the CLAP score, we use a pretrained CLAP model to extract embeddings for the generated audio and its corresponding text prompt, then compute their cosine similarity. This measures the semantic alignment between audio and text. FD compares the distribution of CLAP audio embeddings between real and generated samples by computing the Fréchet Distance between their multivariate Gaussians, capturing semantic feature similarity. FAD compares VGGish [16] embeddings of real and generated audio, reflecting perceptual audio quality. KL divergence is computed between the distributions of CLAP audio embeddings for real and generated samples, quantifying divergence in their semantic representation distributions. Together, these metrics evaluate both the semantic accuracy and perceptual quality of the generated audio. As such, we extract and use only the omnidirectional W channel from our generated audio for these evaluations. When computing the CLAP score, we use the text embedding of the original non-spatial caption. To evaluate the spatial accuracy of generated audio, we utilize the same metrics as the autoencoder, namely, the difference in azimuth, elevation and spatial angle.
III Results and Discussions
Evaluation results of the autoencoder are provided in Table III. STFT and MEL reconstruction results prove that the autoencoder is effective at accurately encoding and decoding spatial audio. Evaluation of the spatial accuracy shows that the autoencoder near perfectly preserves the spatial attributes of the audio during the compression process.
| Model | STFT | MEL | L1() | L1() | |
| SonicMotion | 1.35 | 0.95 | 3.32∘ | 1.43∘ | 3.72∘ |
Evaluation results of our diffusion models are shown in Table II. Comparison against state of the art text-to-audio models such as AudioLDM 2 [17] and Stable Audio Open [18] reveals that both variations of our proposed model retain the ability to generate audio that aligns with the semantics of the sound source. It is important to note that this is strictly a comparison of the plausibility of generated audio and ignores any spatial attributes associated with audio generated from the proposed model. Thus, models are evaluated on audio generated using original non-spatial captions. We make this comparison to show that although our model generates a different format of audio than traditional models, it does not sacrifice any quality in the audio produced.
Results of evaluating our proposed model on its spatial accuracy emphasizes the effectiveness of the additional conditioning parameters in our parametric model as it significantly outperforms its descriptive counterpart. This can be attributed to the wide range of meaning when describing direction with natural language. For example, angles of to the left and to the right can both be reasonably described as “front,” despite being apart. While this error still falls within an acceptable range, the parametric model helps to avoid inherent ambiguity in natural language descriptions.
IV Conclusion
Our proposed SonicMotion, an end-to-end latent diffusion model enables generation of first-order ambisonics audio with precise control over the type of sound, as well as its position and movement in 3D space. We also present a new dataset of spatial audio-caption pairs which includes samples with moving and non-moving sources. Evaluation of our model shows that it is successful in generating spatial audio which aligns with sound and spatial descriptions. Future work may explore extending this implementation to include support for multiple overlapping sources as well as control over the distance of the source with movement such as far to near.